Session: May 1998 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Mantovani at Mon Jun 8 16:27:34 METDST 1998) Success
some interference on bandwith

Onsala (A. Polatidis at Mon Jun 8 16:49:52 METDST 1998) Success

Noto (C.Trigilio at Mon Jun 8 18:08:38 METDST 1998) Success

Jodrell (Alastair Gunn at Tue Jun 9 03:09:44 METDST 1998) Minor failures
Took some time to clear VC alarms at start of experiment. Telescope off-source at about 13:10 due to premature source change by MERLIN, back on source at 13:25. No other known problems.

Cambridge (Alastair Gunn at Tue Jun 9 03:10:10 METDST 1998) Minor failures
Vacuum lost and lost 02:20 scan. Lost vacuum and lost two scans at 04:04 UT, okay for 04:58 UT scan. No other known problems.

Torun (K. Borkowski at Tue Jun 9 08:42:32 METDST 1998) Success
No known problems.

Westerbork (Tony Foley at Wed Jun 10 11:26:34 METDST 1998) Minor failures
Lost first 2 scans as schedule requested observations outside our Hour Angle range. Attenuation increased by 3dB at 06:16UT as VC03 was at saturation. VC05 and VC06 were at the bottom of the band and had no power in the lowest 2MHz.

Calibration (Jean\-Francois Desmurs at Thu Jun 11 12:40:16 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the mai98 directory under the name gp020.antab. Rem: Cm No Tsys Read the comments in the file to get further information.

Effelsberg (dgraham at Fri Jun 12 15:03:26 METDST 1998) Minor failures
source too low until 0140. 1230 antenna tracking problems. Poor recording quality since this should have been on VLBA terminal.